# MotionStream: Real-Time Video Generation with Interactive Motion Controls

### 01 Mar 2026 v3

@@ 1 Introduction
The ultimate goal of motion-controlled video synthesis is to give creators the power of a director’s chair, allowing them to intuitively guide digital actors, objects, and cameras in real time. Although recent video diffusion models have made impressive strides toward this goal (Wang et al., 2024b; Geng et al., 2025; Burgert et al., 2025; Zhang et al., 2025b; Li et al., 2025b; 2024a; Niu et al., 2024; Shi et al., 2024a; Zhou et al., 2025a; Lei et al., 2025), generating high-fidelity videos following user-specified motion trajectories, the experience remains far from interactive.
@@ 2 Related Work
Controllable Video Generation. Enabling precise user control is essential for applying video generative models to diverse downstream applications (Li et al., 2025c; Tu et al., 2025; Bahmani et al., 2024a; Gao et al., 2024c; Wu et al., 2025b; Fu et al., 2025). To this end, a large body of recent research has explored various types of control signals for video generation, such as structure control (Xing et al., 2024; Yang et al., 2025a; Jiang et al., 2025; Pang et al., 2024; Xing et al., 2025), camera control (Gao et al., 2024b; Zheng et al., 2024; He et al., 2024; Bai et al., 2025; Wu et al., 2025a; Yu et al., 2025b; Bahmani et al., 2024b; Yang et al., 2024b; Zheng et al., 2024; 2025), subject control (Huang et al., 2025a; Liu et al., 2025b; Fei et al., 2025; Liu et al., 2025a), and audio control (Tian et al., 2024; Gao et al., 2025; Peng et al., 2024).
As a unique modality that captures underlying video dynamics, motion has become a key conditioning signal for recent video diffusion models. Recent video diffusion models often condition generated videos on diverse forms of motion representations, including optical flow, 2D/3D motion trajectories, bounding boxes, and semantic segmentation (Goldman et al., 2008; Niu et al., 2024; Li et al., 2024b; Wu et al., 2024b; Geng et al., 2025; Zhang et al., 2025b; Gillman et al., 2025; Shi et al., 2024a; Wu et al., 2024a; Gu et al., 2025b; Burgert et al., 2025; Tanveer et al., 2024). Despite their impressive quality, these methods are fundamentally limited to offline processing because they rely on diffusion models with full bidirectional attention, which requires the entire control signal to be known in advance. This constraint prevents their use in real-time, interactive applications.
@@
Another line of research integrates AR and diffusion to enable causal, high-quality video generation (Ruhe et al., 2024; Kim et al., 2024; Xie et al., 2024; Zhang and Agrawala, 2025; Sun et al., 2025; Weng et al., 2024; Liu et al., 2024; Chen et al., 2024; Guo et al., 2025b; Hu et al., 2024; Jin et al., 2025; Gu et al., 2025a; Gao et al., 2024a; Li et al., 2025e; Zhang et al., 2025a). Our work is inspired by the recent paradigm that distills a slow teacher model into a fast AR student for real-time performance (Yin et al., 2025; Huang et al., 2025b; Lin et al., 2025). However, these approaches either exhibit severe color drifts beyond the training horizon or require complex long-video finetuning, which poses challenges for controllable video generation.
@@ 4.2 Quantitative Evaluations
Image Conductor (Li et al., 2025d) AnimateDiff (256P) 2.98 11.30 0.214 0.664 91.64 10.29 0.192 0.644 31.22
@@ Appendix D Training Details
Data Preprocessing. We train our models on two primary data sources: OpenVid-1M (Nan et al., 2024) and synthetic videos generated by larger Wan text-to-video models. For OpenVid-1M, we filter the dataset to 0.6M videos by requiring a minimum of 81 frames and a 16:9 aspect ratio, sampling at 16 FPS. For synthetic data, we use 70K samples for Wan 2.1 (81 frames, 480P resolution, generated by Wan 2.1 14B using text prompts from VidProm (Wang and Yang, 2024)) and 30K publicly available samples for Wan 2.2 (121 frames, 720P resolution, generated by Wan 2.2 5B from the FastVideo team (The FastVideo Team, 2024)).

### 05 Mar 2026 v5

@@ 1 Introduction
The ultimate goal of motion-controlled video synthesis is to give creators the power of a director’s chair, allowing them to intuitively guide digital actors, objects, and cameras in real time. Although recent video diffusion models have made impressive strides toward this goal (Wang et al., 2024b; Geng et al., 2025; Burgert et al., 2025; Zhang et al., 2025b; Li et al., 2025b; c; Niu et al., 2024; Shi et al., 2024a; Zhou et al., 2025a; Lei et al., 2025), generating high-fidelity videos following user-specified motion trajectories, the experience remains far from interactive.
@@ 2 Related Work
Controllable Video Generation. Enabling precise user control is essential for applying video generative models to diverse downstream applications (Li et al., 2025d; Tu et al., 2025; Bahmani et al., 2024a; Gao et al., 2024c; Wu et al., 2025b; Fu et al., 2025). To this end, a large body of recent research has explored various types of control signals for video generation, such as structure control (Xing et al., 2024; Yang et al., 2025a; Jiang et al., 2025; Pang et al., 2024; Xing et al., 2025), camera control (Gao et al., 2024b; Zheng et al., 2024; He et al., 2024; Bai et al., 2025; Wu et al., 2025a; Yu et al., 2025b; Bahmani et al., 2024b; Yang et al., 2024b; Zheng et al., 2024; 2025), subject control (Huang et al., 2025a; Liu et al., 2025b; Fei et al., 2025; Liu et al., 2025a), and audio control (Tian et al., 2024; Gao et al., 2025; Peng et al., 2024).
As a unique modality that captures underlying video dynamics, motion has become a key conditioning signal for recent video diffusion models. Recent video diffusion models often condition generated videos on diverse forms of motion representations, including optical flow, 2D/3D motion trajectories, bounding boxes, and semantic segmentation (Goldman et al., 2008; Niu et al., 2024; Li et al., 2024; Wu et al., 2024b; Geng et al., 2025; Zhang et al., 2025b; Gillman et al., 2025; Shi et al., 2024a; Wu et al., 2024a; Gu et al., 2025b; Burgert et al., 2025; Tanveer et al., 2024). Despite their impressive quality, these methods are fundamentally limited to offline processing because they rely on diffusion models with full bidirectional attention, which requires the entire control signal to be known in advance. This constraint prevents their use in real-time, interactive applications.
@@
Another line of research integrates AR and diffusion to enable causal, high-quality video generation (Ruhe et al., 2024; Kim et al., 2024; Xie et al., 2024; Zhang and Agrawala, 2025; Sun et al., 2025; Weng et al., 2024; Liu et al., 2024; Chen et al., 2024; Guo et al., 2025b; Hu et al., 2024; Jin et al., 2025; Gu et al., 2025a; Gao et al., 2024a; Li et al., 2025f; Zhang et al., 2025a). Our work is inspired by the recent paradigm that distills a slow teacher model into a fast AR student for real-time performance (Yin et al., 2025; Huang et al., 2025b; Lin et al., 2025). However, these approaches either exhibit severe color drifts beyond the training horizon or require complex long-video finetuning, which poses challenges for controllable video generation.
@@ 4.2 Quantitative Evaluations
Image Conductor (Li et al., 2025e) AnimateDiff (256P) 2.98 11.30 0.214 0.664 91.64 10.29 0.192 0.644 31.22
@@ Appendix D Training Details
Data Preprocessing. We train our models on two primary data sources: OpenVid-1M (Nan et al., 2024) and synthetic videos generated by larger Wan text-to-video models. For OpenVid-1M, we filter the dataset to 0.6M videos by requiring a minimum of 81 frames and a 16:9 aspect ratio, sampling at 16 FPS. For synthetic data, we use 70K samples for Wan 2.1 (81 frames, 480P resolution, generated by Wan 2.1 14B using text prompts from VidProm (Wang and Yang, 2024)) and 30K publicly available samples for Wan 2.2 (121 frames, 720P resolution, generated by Wan 2.2 5B from the FastVideo team (The FastVideo Team, )).